Which car has the largest acceleration?

Which car has the largest acceleration?

Goes from 0 m/sm/s to 5.0 m/sm/s in 2.0 ss
Goes from 0 m/sm/s to 3.8 m/sm/s in 1.0 ss
Goes from 0 m/sm/s to 21 m/sm/s in 7.0 ss
Goes from 0 m/sm/s to 7.5 m/sm/s in 5.0 ss
Answer #1

Acceleraion is given by rate of change of velocity i.e a = (Δv) / (Δt) = (v - u)/Δt; where 'v' is the final speed, 'u' is the initial speed and 'Δt' is the time elapsed.

(i) u = 0 m/sec, v = 5 m/sec and  Δt = 2 sec, so acceleration is a = 5/2 = 2.5 m/s2.

(ii)  u = 0 m/sec, v = 3.8 m/sec and  Δt = 1 sec, so acceleration is a = 3.8/1 = 3.8 m/s2.

(iii)  u = 0 m/sec, v = 21 m/sec and  Δt = 7 sec, so acceleration is a = 21/7 = 3 m/s2.

(iv)  u = 0 m/sec, v = 7.5 m/sec and  Δt = 5 sec, so acceleration is a = 7.5/5 = 1.5 m/s2.

So the 2nd car has the largest acceleration that goes from 0 m/s to 3.8 m/s in 1 second.